OpenStudy (anonymous):

what type of right triangle is this if: a=90, b=y+40, c=3y-10

OpenStudy (anonymous):

im assuming a,b,and c are degrees. then b+c=90

OpenStudy (anonymous):

Well, no they are all side measurements

OpenStudy (anonymous):

and also, i am figuring for y.

OpenStudy (anonymous):

is a the hypotenuse?

OpenStudy (anonymous):

no, c is.

OpenStudy (anonymous):

then c^2 - b^2 = a^2

OpenStudy (anonymous):

do that and the quadratic formula to solve for y

OpenStudy (anonymous):

Yes, the pythagorean theorem... and to tell the truth i have absolutely no idea what that means

OpenStudy (anonymous):

what does "that" refer to?

OpenStudy (anonymous):

"that" refers to do that and the quadratic formula to solve for y

OpenStudy (anonymous):

just plug in everything you know into the pythagorean theorem. then rearrange what you get so that you can use the quadratic formula.

OpenStudy (anonymous):

(3y-10)^2 - (y+40)^2 = 90^2

OpenStudy (anonymous):

well.... i did the same thing except i used a squared plus b squared equals c squared

OpenStudy (anonymous):

that works too

OpenStudy (anonymous):

now just rearrange it all so that everything is on one side of the equation

OpenStudy (anonymous):

okay but i dont come up with y being a perfect square.

OpenStudy (anonymous):

i tried this: a=90,b=y+40,c=3y-10 a2+b2=c2 squared=2 so 902+(y+40)2=(3y-10)2 8100+y2+1600=3y2+100 9700+y2=3y2+100 9600+y2=3y2=2y2+100 9600=2y2 4800=y2

OpenStudy (anonymous):

you squared incorrectly. (y+40)^2 is y^2+80y+1600. same with (3y-10)^2, that was squared incorrectly too

OpenStudy (anonymous):

how do you get 80y?

OpenStudy (anonymous):

use the foil method.

OpenStudy (anonymous):

please elaborate....

OpenStudy (anonymous):

hmmm. write out (y+40)x(y+40). then distribute. so it becomes y*y + y*40 + 40*y + 40*40

OpenStudy (anonymous):

ooooohhhhh!!!! okay genius.... (in a good way) thank you so much!

OpenStudy (anonymous):

lol no problem. by the way, the answer isn't a whole number.

OpenStudy (anonymous):

I'm in seventh grade honors and lets just say my teacher isnt exactly very to the point. and oh are you serious? so its not a perfect square?

OpenStudy (anonymous):

nope =(

OpenStudy (anonymous):

ohhh jeez. i was having a hard time the whole time because i thought it HAD to be a perfect square... :) whoops.

OpenStudy (anonymous):

lol.